# Exploring the Challenge and Value of Deep Learning in Automated Skin Disease Diagnosis

### v1

@@ Abstract
Skin cancer is one of the most prevalent and deadly forms of cancer worldwide, which highlights the critical importance of early detection and diagnosis in improving patient outcomes. Deep learning (DL) has shown significant promise in enhancing the accuracy and efficiency of automated skin disease diagnosis, particularly in detecting and evaluating skin lesions and classification. However, there are still several challenges for DL-based skin cancer diagnosis, including complex features, image noise, intra-class variation, inter-class similarity, and data imbalance. By synthesizing recent research, this review discusses innovative approaches to cope with these challenges, such as data augmentation, hybrid models, and feature fusion, etc. Furthermore, the review highlights the integration of DL models into clinical workflows, offering insights into the potential of deep learning to revolutionize skin disease diagnosis and improve clinical decision-making. This article follows a comprehensive methodology based on the PRISMA framework and emphasizes the need for continued advancements to fully unlock the transformative potential of DL in dermatological care.
@@
journal: Neurocomputing
@@ 1 Introduction
As one of the most critical cancers all over the world, skin cancer has an increasing tendency in its incidence over the years [1]. Based on surveys, nearly 108,420 new cases of skin cancer were diagnosed in the United States in 2020 [2], while there are about 97,610 new cases confirmed, with 7,990 deaths estimated in 2022 [3]. Skin cancer is a rapidly growing disease characterized by the abnormal and uncontrolled proliferation of skin cells, primarily caused by prolonged exposure to ultraviolet (UV) radiation from the sun. If left untreated, these cancerous cells may invade deeper tissues and spread to other parts of the body, leading to severe complications. This disease is broadly categorized into two main types that is, melanoma and non-melanoma skin cancer. Melanoma is the most aggressive and life-threatening form of skin cancer. It often appears as a dark, irregularly shaped spot on the skin, which may develop from a pre-existing mole or arise as a new lesion. Non-melanoma skin cancers are more common but generally less aggressive than melanoma. These cancers typically present as persistent lesions, scaly patches, or nodules on sun-exposed areas of the skin [4]. In addition, skin cancer is also typically categorized into two distinct forms: benign and malignant. As a lethal malignancy, skin cancer mainly encompasses 7 types: actinic keratoses (AKIEC), basal cell carcinoma (BCC), benign keratosis-like lesion (BKL), dermatofibroma (DF), melanoma (MEL), melanocytic nevi (NV), and vascular lesion (VASC), all of which can cause serious health risks [5].
@@
[…]
@@ 1.1 Objectives
The objective of this review is to comprehensively analyze the advancements and challenges in the application of deep learning for the automated diagnosis of skin diseases. By evaluating the latest research from 2020 to 2025, the review aims to highlight the effectiveness of various deep learning models in enhancing diagnostic accuracy, overcoming image processing challenges, and addressing data imbalance. Additionally, the review seeks to explore innovative methodologies, including data augmentation, hybrid models, and optimization techniques, that contribute to improving the performance and clinical applicability of deep learning in dermatological diagnostics.
@@ 1.2 Contribution
This review provides a comprehensive overview of the current research status regarding how deep learning models address the 4 challenges mentioned above in the diagnosis of skin diseases, synthesizing findings from various sources to offer a holistic perspective on the field.
By presenting a comparative analysis of different deep learning methodologies, this review offers insights into their respective strengths and weaknesses, assisting researchers and clinicians in selecting the most appropriate approach for specific diagnostic tasks.
@@
The remainder of this paper is organized as follows. Section 2 describes the methodology of this systematic review, including the literature search, inclusion criteria, and study selection process. Section 3 summarizes the preliminary findings derived from the included studies. Section 4 introduces the benchmark datasets and data preprocessing methods used for automated skin disease diagnosis. Section 5 discusses the main challenges faced by current deep learning models—such as image noise, intra-class similarity, and class imbalance—and reviews advanced solutions including data augmentation, transfer learning, and optimized loss functions. Section 6 explores prospective research directions, while Section 7 presents future development recommendations. Then the paper is concluded in this section.
@@ 2 Methodology
To conduct a more comprehensive review, the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework was used for the literature retrieval and selection, as shown in Fig.2. We performed 3 stages for this review. In the identification stage, this study searched the relevant literature through the databases of PubMed, Web of Science, IEEE Xplore, and Elsevier Science Direct, and initially obtained 810 records. Among them, Web of Science and Elsevier Science Direct provided more literature, contributing 314 and 335 articles, respectively, while PubMed and IEEE Xplore contributed 22 and 140 articles, respectively. After removing duplicates, a total of 691 articles were advanced to the next stage. A step-by-step screening strategy was utilized in the screening stage. First, 33 records of ineligible types were excluded, such as reviews, books, patents, conference abstracts, or papers without empirical data. Thus, the number of remaining documents is 658. Subsequently, 515 irrelevant studies were excluded through title and abstract screening, which either did not focus on deep learning or did not address issues related to dermatologic diagnosis, leaving 143 remaining for full-text assessment. In the final stage of eligibility assessment, we further excluded 44 papers with two main problems: first, the studies did not address dermatological diseases, and second, although the studies used deep learning, they did not address the core challenges, such as data imbalance and feature complexity. Then, 98 studies were included in the review analysis. Finally, 91 articles were included after more detailed reading. Zotero is utilized for literature collation and analysis.
@@ 3 Findings
We retrieved publications published from January 2020 to June 2025. It can be learned from Fig. 3 that the publication trend indicates a significant and sustained increase in research activity concerning the application of deep learning to skin disease diagnosis. While the number of publications was relatively modest in 2020 ($n=7$) and 2021 ($n=9$), a marked surge occurred in 2022 with 18 papers published, followed by sustained high outputs in 2023 ($n=20$) and 2024 ($n=22$), suggesting the field entered a phase of accelerated development. This rapid growth may be attributed to advances in deep learning architectures and the availability of high-quality dermoscopic datasets. Although only 15 papers were published in the first half of 2025, this partial data is not indicative of a declining trend but rather reflects the incomplete nature of the annual publication cycle. Overall, the trend reflects increasing academic attention and clinical relevance, with a shift from foundational model development towards addressing key diagnostic challenges, such as feature complexity and data imbalance, and facilitating the translation of AI models into clinical practice.
@@
To address the key points of the included papers, Fig.4 is generated from the article’s title section. The text size is decided by the frequency of words. The high frequency of the terms “skin”, “cancer”, “disease”, and “lesion” indicates that the research mainly focuses on the automatic detection and diagnosis of skin lesions, especially skin cancer. The prominence of keywords such as “classification”, “detection”, and “segmentation” indicates that the research focuses on the classification, detection, and image segmentation of skin lesions to improve diagnostic accuracy. In addition, the words “learningbased”, “hybrid”, “ensemble”, and “transformer” indicate that researchers are exploring different deep learning architectures as well as hybrid models to optimize the diagnostic performance. Meanwhile, the words “imbalance”, “augmentation”, and “explainable” indicate that research is still facing the challenges of data imbalance, feature complexity, and model explainability, and some research attempts to improve the robustness and clinical usability of the models through data augmentation, migration learning, and federated learning. Overall, the research direction in this field covers disease classification, detection optimization, deep learning method improvement, and is constantly exploring stronger models and better data processing solutions to promote the application of intelligent medical image analysis in dermatological diagnosis.
The high-frequency journals retrieved in this study are listed in Table.1. The retrieved journals mainly focus on the high-quality journals of JCR Q1 level, among which IEEE ACCESS has published 7 studies, which indicate the wide application of computer vision and multimedia processing technology in dermatological diagnosis. In addition, journals such as MULTIMEDIA TOOLS AND APPLICATIONS, BIOMEDICAL SIGNAL PROCESSING AND CONTROL, and SCIENTIFIC REPORTS occupy an essential position in the fields of computer-aided diagnosis (CAD), medical image analysis, and clinical medical application, further reflecting the interdisciplinary integration trend in this field. The trend is toward interdisciplinary convergence in the field. The research involves a variety of disciplines such as computer science, engineering, biomedicine, and medicine, indicating that the application of deep learning in dermatologic diagnosis has gradually transformed from methodological research to clinical practice.
Fig. 4: Challenges of current deep learning studies
@@ 4 Research data
High quality dataset is required in the study of skin disease diagnosis based on deep learning. Thus, we summarized benchmark datasets commonly utilized for experimentation, as shown in Table.2.
Table.2 provides an overview of widely used benchmark datasets in automated skin lesion detection research, including Skin Imaging Collaboration(ISIC) Series, HAM10000, PH2, Derm7pt, and SkinCon. These datasets serve as essential resources for developing and evaluating deep learning models in dermatological diagnosis. The ISIC Series, curated by the International Skin Imaging Collaboration, encompasses a diverse range of skin lesion types but is characterized by extreme class imbalance, which poses challenges for model training. Similarly, HAM10000, a large-scale dataset with 10015 dermoscopic images spanning seven common pigmented skin lesion types, also exhibits class imbalance, potentially leading to biased model predictions. In contrast, PH2 is a smaller dataset with 200 images, offering a well-balanced class distribution, making it particularly suitable for model validation and fairness assessment. Additionally, Derm7pt is distinguished by its seven-point checklist annotation system, providing 1011 dermoscopic images with both clinical and imaging data, making it a valuable resource for multimodal deep learning studies. The SkinCon dataset contains 3230 images spanning 48 clinical concepts, enabling fine-grained lesion classification and concept-level labeling. While these datasets significantly contribute to the advancement of deep learning in dermatology, they also present challenges such as class imbalance, small sample sizes, and the need for clinical applicability. Researchers should carefully select appropriate datasets based on their study objectives and employ techniques such as data augmentation, resampling strategies, and multimodal learning to enhance model robustness and improve diagnostic performance in clinical applications.
@@
Dataset Introduction Characteristic
ISIC Series A series of datasets provided by International Skin Imaging Collaboration (ISIC), whose aim is to facilitating research on automated detection of skin lesions. Extremely imbalance; Multiple categories
HAM10000 A dataset, called “Human Against Machine with 10,000 training images”, which contains 10015 dermoscopic images covering seven common pigmented skin lesions. Obvious imbalance; Multiple categories
PH2 This is a dataset of 200 dermoscopic images that were used in a classification study of pigmented skin lesions. Small sample; Balanced data
Derm7pt A dataset labeled based on a 7-point checklist approach, containing 1011 dermoscopic images. Multimodal support; Multiple categories
SkinCon This is a data set of 3230 images, covering 48 clinical concepts. Conceptual level labeling; Multiple categories

### v2

@@ Abstract
Skin cancer is one of the most prevalent and deadly forms of cancer worldwide, highlighting the critical importance of early detection and diagnosis in improving patient outcomes. Deep learning (DL) has shown significant promise in enhancing the accuracy and efficiency of automated skin disease diagnosis, particularly in detecting and classifying skin lesions. However, several challenges remain for DL-based skin cancer diagnosis, including complex features, image noise, intra-class variation, inter-class similarity, and data imbalance. This review synthesizes recent research and discusses innovative approaches to address these challenges, such as data augmentation, hybrid models, and feature fusion. Furthermore, the review highlights the integration of DL models into clinical workflows, offering insights into the potential of deep learning to revolutionize skin disease diagnosis and improve clinical decision-making. This review uniquely integrates a PRISMA-based methodology with a challenge-oriented taxonomy, providing a systematic and transparent synthesis of recent deep learning advances for skin disease diagnosis. It further highlights emerging directions such as hybrid CNN-Transformer architectures and uncertainty-aware models, emphasizing its contribution to future dermatological AI research.
@@
journal: Biomedical Signal Processing and Control
@@ 1 Introduction
As a major global health burden, skin cancer has seen a steady rise in incidence rates in recent years [1]. Based on surveys, nearly 108,420 new cases of skin cancer were diagnosed in the United States in 2020 [2], while there are about 97,610 new cases confirmed, with 7,990 deaths estimated in 2022 [3]. Skin cancer is a rapidly growing disease characterized by the abnormal and uncontrolled proliferation of skin cells, primarily caused by prolonged exposure to ultraviolet (UV) radiation from the sun. If left untreated, these cancerous cells may invade deeper tissues and spread to other parts of the body, leading to severe complications. This disease is broadly categorized into two main types: melanoma and non-melanoma skin cancer. Melanoma is the most aggressive and life-threatening form of skin cancer. It often appears as a dark, irregularly shaped spot on the skin, which may develop from a pre-existing mole or arise as a new lesion. Non-melanoma skin cancers are more common but generally less aggressive than melanoma. These cancers typically present as persistent lesions, scaly patches, or nodules on sun-exposed areas of the skin [4]. In addition, skin cancer is also typically categorized into two distinct forms: benign and malignant. As a lethal malignancy, skin cancer mainly encompasses 7 types: actinic keratoses (AKIEC), basal cell carcinoma (BCC), benign keratosis-like lesion (BKL), dermatofibroma (DF), melanoma (MEL), melanocytic nevi (NV), and vascular lesion (VASC), all of which can cause serious health risks [5].
@@
As a primary technology for the diagnosis of skin cancer, dermoscopy is commonly utilized. However, interpreting dermoscopic images relies heavily on the expertise of professional dermatologists. Although diagnostic accuracy can reach 75%–84% in expert hands, manual assessment is hindered by subjectivity, heavy reliance on experience, and limited reproducibility [19], leading to variability in diagnostic efficiency. Nevertheless, while CAD has the potential to enhance accessibility and accuracy in skin cancer diagnosis [9], numerous technical challenges remain that hinder its widespread effectiveness and reliability. First, uneven lesion color and blurred edges make segmentation and classification difficult. Intra-class variation and inter-class similarity further complicate differentiation, as lesions of the same type can look different, while benign and malignant ones may appear similar. Additionally, image noise and artifacts, such as hair, skin texture, and lighting variations, interfere with accurate detection. The diversity of lesion appearances across different skin tones and demographics adds another layer of complexity. Moreover, the limited generalizability of traditional CAD models, which rely on hand-crafted features and rule-based algorithms, reduces their adaptability to real-world clinical settings. These challenges highlight the need for more advanced deep learning models and improved image processing techniques to enhance CAD accuracy and reliability in skin disease diagnosis.
@@
In contrast to previous reviews, this study not only categorizes recent works under four fundamental challenges but also incorporates a PRISMA-driven selection process to ensure methodological transparency and reproducibility. This dual framework provides both a conceptual and procedural contribution to dermatological AI research.
@@ 1.1 Objectives
This review aims to analyze the advancements and challenges in the application of deep learning to automated skin disease diagnosis. By evaluating the latest research from 2020 to 2025, the review aims to highlight the effectiveness of various deep learning models in enhancing diagnostic accuracy, overcoming image processing challenges, and addressing data imbalance. Additionally, the review seeks to explore innovative methodologies, including data augmentation, hybrid models, and optimization techniques, that contribute to improving the performance and clinical applicability of deep learning in dermatological diagnostics.
@@ 1.2 Contribution
This review provides a comprehensive overview of the current research on how deep learning models address the four challenges mentioned above in automated skin disease diagnosis, synthesizing findings from various sources to offer a holistic perspective on the field. Then, we conduct a comparative analysis of various deep learning methodologies, offering insights into their strengths and weaknesses to guide researchers and clinicians in selecting the most appropriate approach.
Then, a comparative analysis of various deep learning methodologies, is conducted in this paper, offering insights into their strengths and weaknesses to guide researchers and clinicians in selecting the most appropriate approach. Compared with existing surveys, none of them integrates a PRISMA-based methodological rigor with a challenge-oriented taxonomy, nor do they systematically connect technical advances with clinical applicability.
@@
The remainder of this paper is organized as follows. Section 2 details the PRISMA-based methodology and literature search strategy. Section 3 presents a bibliometric analysis of the selected studies, including publication trends and geographical distribution. Section 4 introduces the benchmark datasets and common preprocessing techniques. Section 5 provides a comprehensive review of deep learning methods structured around four key diagnostic challenges: complex features, image noise, intra-class variation, and class imbalance. Section 6 analyzes the dominant technical strategies employed to address these challenges, such as data augmentation and hybrid models. Section 7 discusses frontier advances and emerging paradigms, including multimodal fusion and uncertainty-aware learning. Finally, Section 8 offers a critical discussion on the historical progression and limitations, followed by the conclusion.
@@ 2.1 PRISMA framework
To conduct a more comprehensive review, the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework was used for the literature retrieval and selection, as shown in Fig.2. We performed 3 stages for this review. In the identification stage, this study searched the relevant literature through the databases of PubMed, Web of Science, IEEE Xplore, and Elsevier Science Direct, and initially obtained 810 records. Among them, Web of Science and Elsevier Science Direct provided more literature, contributing 314 and 335 articles, respectively, while PubMed and IEEE Xplore contributed 22 and 140 articles, respectively. After removing duplicates, a total of 691 articles were advanced to the next stage. A step-by-step screening strategy was utilized in the screening stage. First, 33 records of ineligible types were excluded, such as reviews, books, patents, conference abstracts, or papers without empirical data. This resulted in 658 documents. Subsequently, 515 irrelevant studies were excluded through title and abstract screening, as they either did not focus on deep learning or did not address dermatological diagnosis, leaving 143 articles for full-text assessment. In the final stage of eligibility assessment, we further excluded 44 papers with two main problems: first, the studies did not address dermatological diseases, and second, although the studies used deep learning, they did not address the core challenges, such as data imbalance and feature complexity. Initially, 98 studies met the inclusion criteria. Upon a secondary rigorous quality check, 7 additional studies were excluded due to insufficient data or lack of relevance, resulting in a final set of 91 articles included in this review. Zotero is utilized for literature collation and analysis. This structured and reproducible screening process minimizes selection bias and ensures that the included studies form a transparent and methodologically consistent evidence base for the subsequent analysis.
Instead of the usage of standard PRISMA screening procedure, the studies included in this review are organized according to four fundamental challenges frequently encountered in automated dermatological diagnosis, including complex features, image noise, intra and inter-class similarity, and data imbalance. This challenge-oriented structure allows the literature to be examined not merely by model type or algorithmic category like conventional surveys, but by the specific diagnostic obstacles each study seeks to address. Such a mapping facilitates a clearer comparison across heterogeneous methods and highlights how different technical strategies converge on common clinical difficulties. Consequently, this approach facilitates a more coherent synthesis of evidence throughout the review.
2.2 Search strategy
A comprehensive literature search was performed across multiple databases, including PubMed, Web of Science, IEEE Xplore, and Elsevier ScienceDirect. The search was limited to studies published between January 2020 and June 2025 to focus on the most recent advancements in deep learning for automated skin disease diagnosis.
The search utilized the following combination of keywords to ensure a comprehensive retrieval of studies.
(”Deep Learning”) AND (”Dermatology” OR ”Skin Disease” OR ”Skin Cancer”) AND (”Feature Complexity” OR ”Data Imbalance” OR ”Class Imbalance” OR ”Intra-class Variability” OR ”Inter-class Similarity” OR ”Image Artifacts” OR ”Data Augmentation” OR ”Domain Adaptation”)
This keyword strategy was specifically crafted to address key challenges in deep learning-based skin disease diagnosis, including feature complexity, data imbalance, intra-class variability, inter-class similarity, image artifacts, and domain adaptation. The inclusion of terms related to data augmentation also ensured that studies exploring methods to improve model performance were captured.
2.3 Risk of Bias Assessment
To assess the risk of bias in the included studies, we used Newcastle-Ottawa Scale (NOS). Moreover, risk of bias was assessed independently by two reviewers, who evaluated the studies in the following domains: selection bias, performance bias, detection bias, and reporting bias. Any disagreements between reviewers were resolved through discussion, and a third reviewer was consulted when consensus could not be reached.
@@ 3 Findings
We retrieved publications published from January 2020 to June 2025. As illustrated in Fig. 3 that the publication trend indicates a significant and sustained increase in research activity concerning the application of deep learning to skin disease diagnosis. While the number of publications was relatively modest in 2020 ($n=7$) and 2021 ($n=9$), a marked surge occurred in 2022 with 18 papers published, followed by sustained high outputs in 2023 ($n=20$) and 2024 ($n=22$), suggesting the field entered a phase of accelerated development. This rapid growth may be attributed to advances in deep learning architectures and the availability of high-quality dermoscopic datasets. Although only 15 papers were published in the first half of 2025, this partial data is not indicative of a declining trend but rather reflects the incomplete nature of the annual publication cycle. Overall, the trend reflects increasing academic attention and clinical relevance, with a shift from foundational model development towards addressing key diagnostic challenges, such as feature complexity and data imbalance, and facilitating the translation of AI models into clinical practice.
@@
To visualize the thematic focus of the included studies, Fig.4 presents a word cloud generated from the article titles. The font size is proportional to the frequency of the terms. The high frequency of the terms “skin”, “cancer”, “disease”, and “lesion” indicates that the research mainly focuses on the automatic detection and diagnosis of skin lesions, especially skin cancer. The prominence of keywords such as “classification”, “detection”, and “segmentation” indicates that the research focuses on the classification, detection, and image segmentation of skin lesions to improve diagnostic accuracy. In addition, the words “learningbased”, “hybrid”, “ensemble”, and “transformer” indicate that researchers are exploring different deep learning architectures as well as hybrid models to optimize the diagnostic performance. Furthermore, terms such as “imbalance”, “augmentation”, and “explainable” indicate that research is still facing the challenges of data imbalance, feature complexity, and model explainability, and some research attempts to improve the robustness and clinical usability of the models through data augmentation, migration learning, and federated learning. Overall, the research direction in this field covers disease classification, detection optimization, deep learning method improvement, and is constantly exploring stronger models and better data processing solutions to promote the application of intelligent medical image analysis in dermatological diagnosis.
The selected studies were predominantly published in high-impact JCR Q1 journals, as listed in Table 1. Among which, IEEE ACCESS has published 7 studies, which indicate the wide application of computer vision and multimedia processing technology in dermatological diagnosis. In addition, journals such as MULTIMEDIA TOOLS AND APPLICATIONS, BIOMEDICAL SIGNAL PROCESSING AND CONTROL, and SCIENTIFIC REPORTS occupy an essential position in the fields of computer-aided diagnosis (CAD), medical image analysis, and clinical medical application, further reflecting the interdisciplinary integration trend in this field. The trend is toward interdisciplinary convergence in the field. The research involves a variety of disciplines such as computer science, engineering, biomedicine, and medicine, indicating that the application of deep learning in dermatologic diagnosis has gradually transformed from methodological research to clinical practice.
Fig. 4: Word Cloud of the reviewed studies.
@@ 4 Research data
High quality dataset is required in the study of skin disease diagnosis based on deep learning. Thus, we summarized benchmark datasets commonly utilized for experimentation, as shown in Table 2.
Table 2 provides an overview of widely used benchmark datasets in automated skin lesion detection research, including Skin Imaging Collaboration(ISIC) Series, HAM10000, PH2, Derm7pt, and SkinCon. These datasets serve as essential resources for developing and evaluating deep learning models in dermatological diagnosis. The ISIC Series, curated by the International Skin Imaging Collaboration, encompasses a diverse range of skin lesion types but is characterized by extreme class imbalance, which poses challenges for model training. Similarly, HAM10000, a large-scale dataset with 10015 dermoscopic images spanning seven common pigmented skin lesion types, also exhibits class imbalance, potentially leading to biased model predictions. In contrast, PH2 is a smaller dataset with 200 images, offering a well-balanced class distribution, making it particularly suitable for model validation and fairness assessment. Additionally, Derm7pt is distinguished by its seven-point checklist annotation system, providing 1011 dermoscopic images with both clinical and imaging data, making it a valuable resource for multimodal deep learning studies. The SkinCon dataset contains 3230 images spanning 48 clinical concepts, enabling fine-grained lesion classification and concept-level labeling. While these datasets significantly contribute to the advancement of deep learning in dermatology, they also present challenges such as class imbalance, small sample sizes, and the need for clinical applicability. Researchers should carefully select appropriate datasets based on their study objectives and employ techniques such as data augmentation, resampling strategies, and multimodal learning to enhance model robustness and improve diagnostic performance in clinical applications.
@@
Dataset Introduction Characteristic Segmentation Mask Source
ISIC Series Representative example: ISIC-2020 contains 33,126 images with 2 classes (benign 31,926; malignant 581). Earlier years (2016–2017) include more classes and pixel-level lesion masks. Extremely imbalanced; multi-category across years. Available (ISIC-2016/2017); Not available in ISIC-2020. https://challenge.isic-archive.com
HAM10000 Contains 10,015 images across 7 classes (AKIEC, BCC, BKL, DF, MEL, NV, VASC). Widely used for benchmark classification tasks. Obvious imbalance; seven diagnostic categories. Not provided. https://www.kaggle.com/kmader/skin-cancer-mnist-ham10000
PH2 Includes 200 dermoscopic images (common nevi + melanomas) with high-quality lesion annotation. Small dataset; balanced categories. Provided (pixel-level). https://www.fc.up.pt/addi/ph2
Derm7pt Contains 1,011 images labeled according to the 7-point checklist, includes clinical metadata. Multimodal annotation; diagnostic criteria aligned with clinical practice. Not provided. https://derm7pt.org
SkinCon Includes 3,230 clinical images covering 48 dermatological concepts, supporting concept-level modeling. Concept-level labeling; multiple clinical attributes. Not provided. https://skincon-dataset.com
Representative examples of dermoscopic images and their corresponding segmentation masks are shown in Fig.6. The upper row presents the annotated lesion regions, and the lower row shows the original images. These samples illustrate typical variations in lesion color, texture, and boundary clarity, reflecting the visual diversity found in common benchmark datasets.
Fig. 6: Representative samples from major benchmark skin lesion datasets.